### v5

@@ Methods
Quilt-1M [41]. This dataset contains 1 million image-text pairs for histopathology. These data were collected from diverse sources, including medical Twitter, YouTube, and other internet repositories. The dataset comprises over 437,000 unique images paired with 802,000 corresponding text annotations, encompassing a total of 1.469 million UMLS entities extracted from the text. The images cover a range of microscopic magnification scales (0–10x, 10–20x, 20–40x) and were uniformly resized to 299×299 pixels for pretraining.
@@
SICAPv2 [40]. The SICAPv2 dataset containing prostate histology whole slide images with both annotations of global Gleason scores and patch-level Gleason grades. We use the official test set with 2122 images from 31 whole slide images for both annotation-free localization and diagnosis tasks.
@@
WSSS4LUAD [57, 58]. This dataset consists of 67 hematoxylin and eosin stained whole-slide images from Guangdong Provincial People’ Hospital (GDPH) and 20 whole slide images from the cancer genome atlas (TCGA). We filtered out 4693 image patches from the official training split following [45] for our evaluation. Each patch was labeled as tumor epithelial tissue, tumor-associated stroma tissue and normal tissue.
@@
DHMC LUAD [59]. This dataset consists of 143 hematoxylin and eosin stained whole-slide images of lung adenocarcinoma from the Department of Pathology and Laboratory Medicine at Dartmouth-Hitchcock Medical Center (DHMC). All images were labeled into 5 categories, including lepidic, acinar, papillary, micropapillary, and solid. We use the entire dataset for annotation-free diagnosis task following [45].
Approvals for the research were secured from the Institutional Review Board/Ethics Committees of all participating hospitals or institutes. Consent forms were signed by all involved patients. The study was carried out in accordance with the United States Health Insurance Portability and Accountability Act. Furthermore, it was in compliance with the principles of the Declaration of Helsinki, and conformed to policies of the Chinese Center for Disease Control and Prevention, as well as the Chinese Health Laws.
@@
In this section, we introduced the comparative algorithms we employed, which are categorized into four main groups: vision-language pre-training methods, self-supervised pre-training methods, unsupervised anomaly detection methods, and saliency-based methods. Each category will was discussed in detail in the following subsections.
@@
Medical vision-language pre-training Methods Ten medical vision-language pre-training methods were compared in this study, including ConVIRT [60], GLoRIA [30], BioViL [28], and MedKLIP [19], CLIP [42], PLIP [44], BiomedCLIP [43], QuiltNet [41], CONCH [45]. ConVIRT improves visual representations by maximizing the agreement between true image-text pairs versus random pairs via a bidirectional contrastive objective between the image and text modalities. GLoRIA proposes an attention-based framework that learns global and local representations by comparing subregions of images with words in paired reports. BioViL achieves better text modeling by adopting an improved lexicon and innovative language pre-training methods. Then, contrastive learning is utilized with paired text and images. MedKLIP first extracts medical information from radiology reports through a triplet extraction module. Then, it uses a Transformer-based fusion model to combine this information and visual signals at the image block level, enhancing the capability of medical diagnostics. Other models are compared as state-of-the-art algorithms for histopathology, which can be found on Hugging Face: CLIP (https://huggingface.co/openai/clip-vit-base-patch16), PLIP (https://huggingface.co/vinid/plip), BiomedCLIP (https://huggingface.co/microsoft/BiomedCLIP-PubMedBERT_256-vit_base_patch16_224), QuiltNet (https://huggingface.co/wisdomik/QuiltNet-B-32), and CONCH (https://huggingface.co/MahmoodLab/CONCH). These models represent the latest advancements in medical vision-language pre-training. By leveraging large-scale datasets and cutting-edge pre-training techniques, they are capable of more accurate and efficient interpretation of medical images and associated texts, ultimately enhancing diagnostic capabilities.
@@
Self-supervised Pre-training Methods RETFound [13] is a self-supervised pre-training model for retinal images that learns generalizable representations from unlabeled retinal images, providing a basis for label-efficient model adaptation across various applications. In this study, we utilized this model for transfer learning on the retinal fundus dataset using classification annotations and compared it with our proposed annotations-free AFLoc.
@@
Unsupervised Anomaly Detection Methods This study compares ReContrast [24] and RD++ [23], two state-of-the-art unsupervised anomaly detection methods. ReContrast proposes a novel contrastive learning paradigm specifically designed for domain-specific unsupervised anomaly detection. This method enhances the transferability of the pre-trained encoder by jointly optimizing all parameters in an end-to-end fashion. The RD++ approach improves anomaly detection by increasing feature compactness and suppressing anomalous signals through a multi-task learning design. For the feature compactness task, RD++ introduces a self-supervised optimal transport method.
@@
Saliency-based Methods In addition to medical vision-language pre-training methods, 3 comparative saliency methods were also adopted in our study. Grad-CAM [15] uses the gradients flowing into the final convolutional layer for any target category to generate a coarse localization map, highlighting the important areas in the image used for classification. Grad-CAM++ [16] is an extension of Grad-CAM that aims to enhance and refine the visualization of the decision-making process of convolutional neural networks. Compared to the original Grad-CAM approach, Grad-CAM++ can provide clearer and more detailed visual explanations, especially in identifying specific classes in recognition models. Eigen-CAM [17] is a visualization technique to interpret the decisions of convolutional neural networks. It extends the traditional Class Activation Mapping (CAM) method, aiming to offer a more intuitive and detailed visual interpretation. The core idea of Eigen-CAM is to use principal component analysis to extract significant features from the outputs of the convolutional layers.
@@
Disease Localization Methods To evaluate the potential performance improvement of our method with limited location annotations, we compared with the methods proposed by Wang et al. [54] and Li et al. [55]. Wang et al. [54] proposed a weakly-supervised framework for weakly-supervised obect localization by considering large image capacity, various multi-label CNN losses and different pooling strategies. Li et al. [55] proposed a unified approach that effectively leverage class information and limited location annotations.
@@ Acknowledgments
This research was partly supported by the National Natural Science Foundation of China (62222118, U22A2040), Shenzhen Medical Research Fund (B2402047), Shenzhen Science and Technology Program (RCYX20210706092104034, JCYJ20220531100213029), the major key project of Pengcheng Laboratory under grant PCL2023AS1-2, and Key Laboratory for Magnetic Resonance and Multimodality Imaging of Guangdong Province (2023B1212060052).
@@
MIMIC-CXR data is available at https://physionet.org/content/mimic-cxr/2.0.0 for users with credentialed access. RSNA Pneumonia data is available at https://www.rsna.org/rsnai/ai-image-challenge/RSNA-Pneumonia-Detection-Challenge-2018. MS-CXR data is available at https://physionet.org/content/ms-cxr/0.1/. CheXlocalize data is available at https://stanfordaimi.azurewebsites.net/datasets/abfb76e5-70d5-4315-badc-c94dd82e3d6d. COVID Rural data is available at https://www.cancerimagingarchive.net/collection/covid-19-ar/. SIIM-ACR Pneumothorax data is available at https://www.kaggle.com/c/siim-acr-pneumothorax-segmentation. NIH ChestXray4 data is available at https://nihcc.app.box.com/v/ChestXray-NIHCC. ChestX-Det10 data is available at https://github.com/Deepwise-AILab/ChestX-Det10-Dataset. Quilt-1M data is available at https://github.com/wisdomikezogwo/quilt1m. SICAPv2 data is available at https://data.mendeley.com/datasets/9xxm58dvs3/2. WSSS4LUAD data is available at https://wsss4luad.grand-challenge.org/. DHMC LUAD data is available at https://bmirds.github.io/LungCancer/.
@@
a
b
| Methods | RSNA Pneumonia | MS-CXR | Covid-Rural |
@@
| Methods | Choroidal neovascularization | Drusen | Intraretinal hemorrhage | Mean |
| IoU | ||||
| ReContrast | 0.080 (0.068, 0.093) | 0.027 (0.019, 0.035) | 0.035 (0.031, 0.038) | 0.047 (0.039, 0.055) |
| RD++ | 0.121 (0.112, 0.131) | 0.062 (0.051, 0.077) | 0.051 (0.047, 0.054) | 0.078 (0.070, 0.087) |
| ConVIRT | 0.131 (0.119, 0.144) | 0.120 (0.099, 0.140) | 0.079 (0.074, 0.085) | 0.110 (0.097, 0.123) |
| GLoRIA | 0.143 (0.130, 0.156) | 0.107 (0.085, 0.128) | 0.075 (0.070, 0.080) | 0.108 (0.095, 0.121) |
| AFLoc | 0.321 (0.303, 0.337) | 0.147 (0.125, 0.170) | 0.097 (0.091, 0.102) | 0.188 (0.173, 0.203) |
| Dice | ||||
| ReContrast | 0.124 (0.107, 0.141) | 0.046 (0.034, 0.061) | 0.064 (0.058, 0.069) | 0.078 (0.066, 0.090) |
| RD++ | 0.194 (0.179, 0.209) | 0.105 (0.087, 0.124) | 0.092 (0.087, 0.098) | 0.130 (0.118, 0.144) |
| ConVIRT | 0.205 (0.187, 0.223) | 0.187 (0.156, 0.215) | 0.141 (0.132, 0.149) | 0.178 (0.158, 0.199) |
| GLoRIA | 0.224 (0.205, 0.242) | 0.168 (0.137, 0.197) | 0.134 (0.126, 0.142) | 0.175 (0.156, 0.194) |
| AFLoc | 0.451 (0.430, 0.470) | 0.225 (0.192, 0.257) | 0.170 (0.160, 0.178) | 0.282 (0.261, 0.302) |
| CNR | ||||
| ReContrast | 0.901 (0.850, 0.954) | 0.421 (0.329, 0.510) | 0.253 (0.231, 0.274) | 0.525 (0.470, 0.579) |
| RD++ | 0.673 (0.635, 0.707) | 0.457 (0.395, 0.521) | 0.210 (0.191, 0.229) | 0.447 (0.407, 0.486) |
| ConVIRT | 0.359 (0.278, 0.447) | 0.652 (0.517, 0.778) | 0.255 (0.231, 0.278) | 0.422 (0.342, 0.501) |
| GLoRIA | 0.493 (0.427, 0.563) | 0.655 (0.512, 0.787) | 0.216 (0.194, 0.237) | 0.455 (0.378, 0.529) |
| AFLoc | 1.452 (1.368, 1.534) | 0.925 (0.783, 1.063) | 0.372 (0.349, 0.396) | 0.916 (0.833, 0.998) |
@@
| Disease | ReContrast | RD++ | GLoRIA | BioViL | MedKLIP | AFLoc |
| Pneumonia | 0.090 (0.080, 0.100) | 0.093 (0.084, 0.103) | 0.290 (0.265, 0.314) | 0.305 (0.284, 0.328) | 0.296 (0.277, 0.315) | 0.412 (0.386, 0.435) |
| Atelectasis | 0.079 (0.064, 0.095) | 0.075 (0.063, 0.088) | 0.303 (0.266, 0.339) | 0.266 (0.222, 0.312) | 0.324 (0.294, 0.355) | 0.381 (0.344, 0.417) |
| Consolidation | 0.098 (0.082, 0.115) | 0.120 (0.101, 0.139) | 0.304 (0.277, 0.330) | 0.299 (0.273, 0.325) | 0.265 (0.238, 0.289) | 0.371 (0.345, 0.395) |
| Pneumothorax | 0.053 (0.045, 0.062) | 0.093 (0.084, 0.103) | 0.116 (0.099, 0.133) | 0.109 (0.094, 0.124) | 0.092 (0.078, 0.106) | 0.117 (0.099, 0.135) |
| Pleural Effusion | 0.061 (0.048, 0.074) | 0.048 (0.037, 0.058) | 0.330 (0.304, 0.357) | 0.227 (0.205, 0.250) | 0.216 (0.195, 0.237) | 0.350 (0.326, 0.375) |
| Lung Opacity | 0.061 (0.046, 0.078) | 0.078 (0.060, 0.098) | 0.196 (0.164, 0.230) | 0.178 (0.148, 0.209) | 0.197 (0.167, 0.228) | 0.327 (0.290, 0.364) |
| Cardiomegaly | 0.131 (0.121, 0.140) | 0.073 (0.066, 0.080) | 0.408 (0.398, 0.420) | 0.267 (0.247, 0.286) | 0.394 (0.384, 0.404) | 0.367 (0.357, 0.377) |
| Edema | 0.120 (0.092, 0.150) | 0.162 (0.129, 0.193) | 0.201 (0.165, 0.236) | 0.171 (0.135, 0.206) | 0.326 (0.285, 0.366) | 0.269 (0.237, 0.299) |
@@
| Hernia | 0.485 | 0.417 | 0.494 | 0.450 | 0.698 | 0.783 | 0.841 | 0.847 (0.798, 0.884) |
@@
| AFLoc | 0.497 (0.471, 0.523) | 0.679 (0.663, 0.693) | 0.504 (0.371, 0.703) |
@@
| Different Levels of Features | RSNA Pneumonia | COVID Rural | MS-CXR | CheXlocalize | Mean | ||
| Word – Shallow local | Sentence – Deep local | Report – Global | |||||
| ✓ | 0.239 | 0.147 | 0.180 | 0.133 | 0.175 | ||
| ✓ | 0.281 | 0.169 | 0.318 | 0.301 | 0.267 | ||
| ✓ | 0.191 | 0.056 | 0.105 | 0.186 | 0.135 | ||
| ✓ | ✓ | 0.329 | 0.177 | 0.264 | 0.285 | 0.264 | |
| ✓ | ✓ | 0.254 | 0.159 | 0.175 | 0.200 | 0.197 | |
| ✓ | ✓ | 0.285 | 0.211 | 0.325 | 0.305 | 0.282 | |
| ✓ | ✓ | ✓ | 0.342 | 0.212 | 0.324 | 0.318 | 0.299 |
@@
| BioClinicalBERT | 0.324 | 0.461 | 1.634 |

### v7

@@ Methods
Quilt-1M [quilt]. This dataset contains 1 million image-text pairs for histopathology. These data were collected from diverse sources, including medical Twitter, YouTube, and other internet repositories. The dataset comprises over 437,000 unique images paired with 802,000 corresponding text annotations, encompassing a total of 1.469 million UMLS entities extracted from the text.
@@
SICAPv2 [sicap]. The SICAPv2 dataset containing prostate histology whole slide images with both annotations of global Gleason scores and patch-level Gleason grades. We use the official test set with 2,122 images from 31 whole slide images for both annotation-free localization and diagnosis tasks.
@@
WSSS4LUAD [wsss4luad, wsss4luad2]. This dataset consists of 67 hematoxylin and eosin stained whole-slide images from Guangdong Provincial People’s Hospital (GDPH) and 20 whole slide images from the cancer genome atlas (TCGA). We filtered out 4,693 image patches from the official training split following [conch] for our evaluation. Each patch was labeled as tumor epithelial tissue, tumor-associated stroma tissue and normal tissue.
@@
DHMC LUAD [dhmcluad]. This dataset consists of 143 hematoxylin and eosin stained whole-slide images of lung adenocarcinoma from the Department of Pathology and Laboratory Medicine at Dartmouth-Hitchcock Medical Center (DHMC). All images were labeled into 5 categories, including lepidic, acinar, papillary, micropapillary, and solid. We use the entire dataset for annotation-free diagnosis task following [conch].
@@
In this section, we introduced the comparative algorithms we employed, which are categorized into four main groups: vision-language pre-training methods, self-supervised pre-training methods, unsupervised anomaly detection methods, and saliency-based methods. Each category will be discussed in detail in the following subsections.
@@
Medical Vision-Language Pre-Training Methods Ten medical vision-language pre-training methods were compared in this study, including ConVIRT [convirt], GLoRIA [huang2021gloria], BioViL [boecking2022making], MedKLIP [wu2023medklip], CLIP [clip], PLIP [plip], BiomedCLIP [biomedclip], QuiltNet [quilt], and CONCH [conch]. ConVIRT improves visual representations by maximizing the agreement between true image-text pairs versus random pairs via a bidirectional contrastive objective between the image and text modalities. GLoRIA proposes an attention-based framework that learns global and local representations by comparing subregions of images with words in paired reports. BioViL achieves better text modeling by adopting an improved lexicon and innovative language pre-training methods. MedKLIP extracts medical information from radiology reports via a triplet extraction module and uses a transformer-based fusion model to integrate this information with image features. Other models are compared as state-of-the-art algorithms for histopathology, which can be found on Hugging Face. These models represent the latest advancements in medical vision-language pre-training.
@@
Self-Supervised Pre-Training Methods RETFound [zhou2023foundation] is a self-supervised pre-training model for retinal images that learns generalizable representations from unlabeled retinal images, providing a basis for label-efficient model adaptation across various applications. In this study, we used this model for transfer learning on the retinal fundus dataset using classification annotations and compared it with our proposed annotations-free AFLoc.
@@
Unsupervised Anomaly Detection Methods This study compares ReContrast [ReContrast] and RD++ [rd++], two state-of-the-art unsupervised anomaly detection methods. ReContrast proposes a novel contrastive learning paradigm specifically designed for domain-specific unsupervised anomaly detection. This method enhances the transferability of the pre-trained encoder by jointly optimizing all parameters in an end-to-end fashion. The RD++ approach improves anomaly detection by increasing feature compactness and suppressing anomalous signals through a multi-task learning design.
@@
Saliency-Based Methods In addition to medical vision-language pre-training methods, 3 comparative saliency methods were also adopted in our study. Grad-CAM [gradcam] uses the gradients flowing into the final convolutional layer for any target category to generate a coarse localization map, highlighting the important areas in the image used for classification. Grad-CAM++ [gradcam++] is an extension of Grad-CAM that aims to enhance and refine the visualization of the decision-making process of convolutional neural networks. Compared to the original Grad-CAM approach, Grad-CAM++ can provide clearer and more detailed visual explanations, especially in identifying specific classes in recognition models. Eigen-CAM [eigencam] is a visualization technique to interpret the decisions of convolutional neural networks. It extends the traditional Class Activation Mapping (CAM) method, aiming to offer a more intuitive and detailed visual interpretation.
@@
Disease Localization Methods To evaluate the potential performance improvement of our method with limited location annotations, we compared with the methods proposed by Wang et al. [wang_chestx-ray8_2017] and Li et al. [li2018thoracic]. Wang et al. [wang_chestx-ray8_2017] proposed a weakly-supervised framework for weakly-supervised object localization by considering large image capacity, various multi-label CNN losses and different pooling strategies. Li et al. [li2018thoracic] proposed a unified approach that effectively leverage class information and limited location annotations.
@@ Acknowledgments
This research was partly supported by the National Natural Science Foundation of China (62222118 to S.S.W., U22A2040 to S.S.W., W2431057 to K.Z.); Shenzhen Medical Research Fund (B2402047 to S.S.W.); the major key project of Pengcheng Laboratory under grant PCL2024A06 (to H.Y., J.L., and W.H.); Key Laboratory for Magnetic Resonance and Multimodality Imaging of Guangdong Province (2023B1212060052 to S.S.W.); Youth Innovation Promotion Association CAS (to S.S.W.); the Macau Science and Technology Development Fund, Macao (0007/2020/AFJ to K.Z., 0070/2020/A2 to K.Z. and 0003/2021/AKP to K.Z.); Macau University of Science and Technology; State Key Laboratory for Eye Health, Eye Hospital and Institute for Advanced Study on Eye Health and Diseases, Wenzhou Medical University.
Author contributions
S.S.W. and H.Y. conceived the idea. S.S.W., H.Y., H.-Y.Z., and K.Z. designed the experiments. H.Y., J.L., and W.H. implemented and conducted the experiments. H.Y., H.-Y.Z., J.L., W.H., and C.L. analyzed the data and experimental results. S.S.W., H.Y., H.-Y.Z., J.L., W.H., and C.L. wrote the paper. K.Z., Z.L., and Y.G. contributed the retinal fundus dataset and analysis tools. Q.L., Y.L., Q.Y., S.W., T.T., and H.Z. participated in discussions and contributed critical feedback. All authors reviewed and approved the final version of the paper.
@@
MIMIC-CXR data is available at https://physionet.org/content/mimic-cxr/2.0.0 for users with credentialed access. RSNA Pneumonia data is available at https://www.rsna.org/rsnai/ai-image-challenge/RSNA-Pneumonia-Detection-Challenge-2018. MS-CXR data is available at https://physionet.org/content/ms-cxr/0.1/. CheXlocalize data is available at https://stanfordaimi.azurewebsites.net/datasets/23c56a0d-15de-405b-87c8-99c30138950c. COVID Rural data is available at https://www.cancerimagingarchive.net/collection/covid-19-ar/. SIIM-ACR Pneumothorax data is available at https://www.kaggle.com/c/siim-acr-pneumothorax-segmentation. NIH ChestXray4 data is available at https://nihcc.app.box.com/v/ChestXray-NIHCC. ChestX-Det10 data is available at https://github.com/Deepwise-AILab/ChestX-Det10-Dataset. Quilt-1M data is available at https://github.com/wisdomikezogwo/quilt1m. SICAPv2 data is available at https://data.mendeley.com/datasets/9xxm58dvs3/2. WSSS4LUAD data is available at https://wsss4luad.grand-challenge.org/. DHMC LUAD data is available at https://bmirds.github.io/LungCancer/. CXR-LT data is available at https://physionet.org/content/cxr-lt-iccv-workshop-cvamd/2.0.0/. The Retinal Fundus data were collected by the hospitals. Owing to patient-privacy considerations, they are not publicly available. Researchers may contact the corresponding authors to discuss data availability for research purposes upon reasonable request.
@@
Supplementary Tables
| Methods | RSNA Pneumonia | MS-CXR | COVID Rural |
@@
Bold values represent the highest performance score among the compared methods.
| Methods |
| IoU | ||||||||||
| ReContrast |
| RD++ |
| ConVIRT |
| GLoRIA |
| AFLoc |
| Dice | ||||||||||
| ReContrast |
| RD++ |
| ConVIRT |
| GLoRIA |
| AFLoc |
| CNR | ||||||||||
| ReContrast |
| RD++ |
| ConVIRT |
| GLoRIA |
| AFLoc |
Bold values represent the highest performance score among the compared methods.
@@
Bold values represent the highest performance score among the compared methods.
| Disease | ReContrast | RD++ | GLoRIA | BioViL | MedKLIP | AFLoc | ||||||||||||
| Pneumonia |
| Atelectasis |
| Consolidation |
| Pneumothorax |
| Pleural Effusion |
| Lung Opacity |
| Cardiomegaly |
| Edema |
Bold values represent the highest performance score among the compared methods.
@@
| Hernia | 0.485 | 0.417 | 0.494 | 0.450 | 0.698 | 0.783 | 0.841 | 0.847 (0.798, 0.884) | |
| CXR-LT | Task 2 | 0.526 | 0.500 | 0.643 | 0.650 | 0.648 | 0.669 | 0.677 | 0.735 (0.715, 0.756) |
| Task 3 | 0.560 | 0.504 | 0.650 | 0.611 | 0.585 | 0.638 | 0.588 | 0.726 (0.716, 0.736) |
Bold values represent the highest performance score among the compared methods.
@@
| AFLoc | 0.512 (0.490, 0.533) | 0.704 (0.692, 0.717) | 0.442 (0.325, 0.521) |
Bold values represent the highest performance score among the compared methods.
@@
Bold values represent the highest performance score among the compared methods.
@@
Bold values represent the highest performance score among the compared methods.
| Different Levels of Features |
MS-CXR | CheXlocalize | Mean | ||||||
| ✓ | 0.239 | 0.147 | 0.180 | 0.133 | 0.138 | ||||||
| ✓ | 0.281 | 0.169 | 0.318 | 0.301 | 0.272 | ||||||
| ✓ | 0.191 | 0.056 | 0.105 | 0.186 | 0.145 | ||||||
| ✓ | ✓ | 0.329 | 0.177 | 0.264 | 0.285 | 0.271 | |||||
| ✓ | ✓ | 0.254 | 0.159 | 0.175 | 0.200 | 0.179 | |||||
| ✓ | ✓ | 0.285 | 0.211 | 0.325 | 0.305 | 0.281 | |||||
| ✓ | ✓ | ✓ | 0.342 | 0.211 | 0.324 | 0.318 | 0.299 | ||||
Bold values represent the highest performance score among the compared methods.
@@
| BioClinicalBERT | 0.324 | 0.462 | 1.636 |
Bold values represent the highest performance score among the compared methods.